The boat is heading due north as it crosses a wide river with a velocity of 10 km/h relative to water. The river has a maintain velocity of 5 km/h due east. Determine the velocity of the boat with respect to an observer on the riverbank.


1
Expert's answer
2021-07-13T10:42:02-0400
"v=\sqrt{10^2+5^2}=11.2\frac{m}{s}\\
\theta=\arctan{\frac{5}{10}}=26.6\degree\text{ E of N}"